MrMudd183:

Can someone help me on this geometry problem please? *I will post the snippet on what I need help on*

MrMudd183:

1 attachment
jhonyy9:

zyx is a right triangle ?

MrMudd183:

Yea

jhonyy9:

the law of cos do you know in a right triangle ?

MrMudd183:

Yea. Isn't it adjacent over hypotenuse?

jhonyy9:

sure ,this is

jhonyy9:

so than cos x = ?

MrMudd183:

yea

jhonyy9:

yea what ? cos x = ?

MrMudd183:

oop, I thought you was asking if I needed to find the cos of x. I think it's 9/12. Do I divide that then find the inverse cos of that to get x?

jhonyy9:

no you ve wrote above the law of cos in a right triangle - so just you need use it and get the cos x value correct

MrMudd183:

I took 9/15 and did the inverse cos and I got 53.1301

jhonyy9:

where you see there 9 i see there 10 - 12 and 15

MrMudd183:

Here's a better image of the triangle. I was trying to snip it without getting the question number in it. This is for a study guide

1 attachment
jhonyy9:

ok so than the cos x = 9/15

MrMudd183:

I would then divide that and take the inverse cos to get the answer? If my thinking is correct, would I get 53.1301?

jhonyy9:

what you have with this inverse ? i dont see this never you need getting the value of cos zxy = ?

MrMudd183:

I took the cos of 9/15 and got 0.9999

jhonyy9:

how you get this 0,9 9/15 = 3/5 = ?

jhonyy9:

so just in your had 3/5 = 0, 30/5 = 6 so 3/5 = 0,6 ok ?

jhonyy9:

so simply and easy

MrMudd183:

Thank ya

jhonyy9:

yw and anytime just tag me than i m online and i will try help you always

MrMudd183:

I'll keep that in mind :)

jhonyy9:

ok - bc. is my pleasure to help in math always - i relax with ...